# MeSS: City Mesh-Guided Outdoor Scene Generation with Cross-View Consistent Diffusion

### v1

@@ 3.2 Stage I: Gaussian Field Construction from Key Views
Two multi-conditioned ControlNets (Zhang, Rao, and Agrawala 2023) are introduced to generate $N$ key views in autoregressive fashion, moving backwards from the end to the start of the camera path [^1] Views are indexed in driving direction.. First, ControlNet-S takes control signals $\{\boldsymbol{d}^{N},\boldsymbol{s}^{N},\boldsymbol{n}^{N}\}$ (depth, semantics and normals) at camera pose $\mathcal{C}^{N}$ to generate view image $\boldsymbol{I}^{N}$. As single-channel disparity maps may lose the ability to discriminate depth values in the far field, so we encode depth as a colormap. This ensures depth variations are retained both near the camera and in the far field.
@@ 3.3 Stage II: Gaussian Field Densification
Specifically, AGInpaint rectifies the prediction of the score function ${\boldsymbol{\epsilon}}_{\theta}$ to the direction, that minimizes the discrepancy between diffusion prediction ${\tilde{\mathbf{x}}}_{\text{0}}$  [^2] By a slight abuse of notation, we skip the step from pixel space to latent space for simplicity. and known RGB value $\mathbf{x}_{\text{0}}$ in the unmasked region $\mathcal{M}$. In simple words, the diffusion process should predict correct RGB values for known regions. This objective guides the diffusion model to predict aligned inpainting with neighbor pixels. At each sampling step, the rectification process is repeated $N_{g}$ times, incrementally updating the latent with a low learning rate $lr$. This corresponds to a gradient descent towarda a local optimum, thus preserving sharper details and more consistent texture within the reference image. The guided rectification is beneficial already in the initial stages of the sampling process: at earlier timesteps it aligns the overall color palette and dominant hues with those of the reference image, at later timesteps, the progressive refinement corrects fine-grained details and textures. For optimal performance of the guided inapinting we adopt LCM. Its self-consistency property Eq. 1 enables consistent estimation of the original sample $\mathbf{x}_{0}$ and stabilizes the guidance across different timesteps during diffusion sampling.
@@ 4.2 Comparison with other methods
For StreetScapes  (Deng et al. 2024), we directly use the results reported in their paper since their code is not publicly available. However, forming a fair comparison is challenging, as their model is trained on a significantly larger dataset compared to ours [^3] Even for their experiment on the London patch, they utilize approximately 20 times more data than we do.. Given that our generated sequence is longer than theirs (200 frames vs. 64 frames), we evaluate their results within the range of 32 to 64 frames. In terms of quality, our results are on par with theirs, while we outperform them in terms of temporal consistency. This advantage can be attributed to our method’s effective utilization of geometric priors, which plays a key role in achieving superior consistency.

### v2

@@ 3.2 Stage I: Gaussian Field Construction from Key Views
Two multi-conditioned ControlNets (Zhang, Rao, and Agrawala 2023) are introduced to generate $N$ key views in autoregressive fashion, moving backwards from the end to the start of the camera path***Views are indexed in driving direction.. First, ControlNet-S takes control signals $\{\boldsymbol{d}^{N},\boldsymbol{s}^{N},\boldsymbol{n}^{N}\}$ (depth, semantics and normals) at camera pose $\mathcal{C}^{N}$ to generate view image $\boldsymbol{I}^{N}$. As single-channel disparity maps may lose the ability to discriminate depth values in the far field, so we encode depth as a colormap. This ensures depth variations are retained both near the camera and in the far field.
@@ 3.3 Stage II: Gaussian Field Densification
Specifically, AGInpaint rectifies the prediction of the score function ${\boldsymbol{\epsilon}}_{\theta}$ to the direction, that minimizes the discrepancy between diffusion prediction ${\tilde{\mathbf{x}}}_{\text{0}}$ By a slight abuse of notation, we skip the step from pixel space to latent space for simplicity. and known RGB value $\mathbf{x}_{\text{0}}$ in the unmasked region $\mathcal{M}$. In simple words, the diffusion process should predict correct RGB values for known regions. This objective guides the diffusion model to predict aligned inpainting with neighbor pixels. At each sampling step, the rectification process is repeated $N_{g}$ times, incrementally updating the latent with a low learning rate $lr$. This corresponds to a gradient descent towarda a local optimum, thus preserving sharper details and more consistent texture within the reference image. The guided rectification is beneficial already in the initial stages of the sampling process: at earlier timesteps it aligns the overall color palette and dominant hues with those of the reference image, at later timesteps, the progressive refinement corrects fine-grained details and textures. For optimal performance of the guided inapinting we adopt LCM. Its self-consistency property Eq. 1 enables consistent estimation of the original sample $\mathbf{x}_{0}$ and stabilizes the guidance across different timesteps during diffusion sampling.
@@ 4.2 Comparison with other methods
For StreetScapes  (Deng et al. 2024), we directly use the results reported in their paper since their code is not publicly available. However, forming a fair comparison is challenging, as their model is trained on a significantly larger dataset compared to oursEven for their experiment on the London patch, they utilize approximately 20 times more data than we do.. Given that our generated sequence is longer than theirs (200 frames vs. 64 frames), we evaluate their results within the range of 32 to 64 frames. In terms of quality, our results are on par with theirs, while we outperform them in terms of temporal consistency. This advantage can be attributed to our method’s effective utilization of geometric priors, which plays a key role in achieving superior consistency.